# OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows

### v1

@@ 4.1 Evolution of H ii Regions
H ii regions are dynamic environments with radiation and winds from O and B stars shaping and enriching the proximal ISM. Winds launched from stars carry out recently formed metals dredged up to the surface and subsequently sweep up surrounding material. Radiation heats gas, often through photoionization, and imparts momentum to the ISM, which can result in additional launching or driving of winds, bubbles, and outflowing gas (e.g., Castor1975b; Weaver1977; Dyson1977; Dyson1979; Oey1997; Dopita2005). While stellar population models often contain predictions for these feedback effects (e.g., 1999ApJS..123....3L; starburst99; Eldrige2017; Byrne2022; Hawcroft2025), empirically constraining the kinematics and related mechanical feedback in H ii regions remains an active area of research (e.g., models and simulations by Geen2020; Geen2023; Lancaster2025a; Lancaster2025b) with implications for understanding chemical enrichment (e.g., Esteban1995; LopezSanchez2007; Stock2011) and dispersal timescales (e.g., Dopita2005; Menon2024).
One such kinematically rich and complex system is the giant extragalactic H ii region NGC 5471 located in M101. Many knots within this region exhibit broad and sometimes asymmetric nebular lines which cannot be described by a single Gaussian (e.g., Chu1986; Bresolin2020). To investigate the evolution of H ii regions in this case-study, we apply OutLines to the [S ii] $\lambda\lambda 6716,6731$ doublet, which traces the density and kinematics of low ionization gas, in Knot A of NGC 5471, the brightest knot identified by Skillman1985, as observed with the $R=6000$ MEGARA IFU spectrograph on the Gran Telescopia Canarias (Arellano-Cordova in prep, PI Esteban). For our modeling, we use a 2.5′′ aperture extraction of the Knot A spectrum from the IFU cube. We obtain atomic data from NIST (NIST_ASD) for wavelengths (Kaufman1993) and spontaneous emission coefficients (transition probabilities) (Podobedova2009). We use collision strengths from Tayal2010. First, we fit the adjacent featureless continuum with a third-order polynomial, an appropriate approximation in the case of extended H ii regions which do not suffer appreciably from contamination such as stellar absorption except near permitted lines from, e.g., H i and He i. We subtract this continuum before modeling the line profile with OutLines with the default options. Then, we sample the posterior on the OutLines profile parameters to constrain the outflow and static gas properties. No LSF corrections are necessary as $R=6000$ is sufficiently high to constrain the kinematics (see §3.2). We compare the best-fit model to the observed line profiles in Figure 16.
@@
From the best-fit line fluxes, we determine the electron density of both the nebula and outflow components from the [S ii] 6716/6731{6716}/{6731} flux ratio using PyNeb (2015A&A...573A..42L). The static gas flux ratio of 1.459 is in the low-density limit, requiring $n_{e}<10\rm\ cm^{-3}$. For the outflowing gas component, the doublet ratio is 0.831, which is where [S ii] is most sensitive simultaneous to density and to temperature. For a range in $T_{e}$ from 10 to 15 $\rm kK$, the inferred outflow density is anywhere from $970$ to $1140\rm~cm^{-3}$, with density values increasing with the assumed temperature. The implication from these densities is that the outflow is roughly two orders of magnitude more dense than the gas within the static nebular gas. While this increase in density may be due to a sweeping up of material in the wind, this dramatic change in density coupled with the supersonic flow could instead arise from compression by shocked gas as it rapidly cools via radiation (e.g., Flury2025b).
The slope of the velocity field, $\beta=4$, corresponds to gradual acceleration likely caused by optically thick radiation pressure from early type stars (e.g., 1977ApJ...213..737B). While Bresolin2020 argue that the non-thermal radio SEDs of Sramek1986 are clear-cut evidence for supernova-driven bubbles, the line profile shape indicates a continuous wind rather than shells or bubbles and, from modeling with OutLines, disfavors the supernova scenario. Indeed, the OutLines results are very much in agreement with the fact that Knot A features the a very young ($3\rm~Myr$) cluster–in fact the youngest in the entirety of NGC 5471 (e.g., GarciaBenito2011, based on HST photometric color-magnitude diagrams). Knot A is also very massive ($10^{4}\rm~M_{\odot}$) and requires a significant population of O stars to produce the $Q=4\times 10^{51}\rm~phot~s^{-1}$ associated with the star cluster (GarciaBenito2011). The age and mass of the stellar population indicate sufficient radiation available in the UV to radiatively drive a supersonic wind. Shocks triggered by this wind could explain the non-thermal emission in the radio in addition to the dense gas. Inhomogeneities in the gas distribution can readily dissipate the outward force due to pressure from the hot shocked gas (e.g., Dyson1977), allowing radiation pressure to dominate the acceleration of material out from the H ii region. Our results indicating escape of hot gas are consistent with the HarperClark2009 model for the Carina Nebula and the Lopez2011 multiwavelength study of the massive extragalactic H ii region 30 Doradus, both of which found strong evidence for radiation pressure driving and a deficit of X-ray flux associated with shocked gas.
@@ 4.2 Lyman Continuum Escape from a Super Star Cluster
Starburst galaxy J1044+0353 is a low mass ($M_{\star}\approx 10^{6.8}\rm~M_{\odot}$) Lyman continuum emitter candidate with $f_{esc}^{\rm Lyc}$ as high as 60% (Parker2025) and, as a nearby ($z=0.01287$) Green Pea, is a local analog for star-forming galaxies at cosmic dawn that are actively involved in re-ionizing the Universe (Martin2024). Recent IFU observations of J1044+0353 reveal expanding shells of gas at the periphery of the nebular gas flux distribution (Martin2024), indicative of blowout due to SNe from an aging stellar population 10-20 Myr after a burst of star formation (Peng2023). However, the central core of the galaxy hosts a super star cluster with several young (<2<2 Myr) stellar populations which may be responsible for driving additional feedback on shorter timescales (Olivier2022), providing a unique laboratory to explore the impact of young stars on their immediate environment in the context of stochastic, or bursty, star formation using OutLines.
@@
To undertake this investigation, we draw on archival HST/COS G130M FUV spectroscopy (R=10  [^0] R=10\ 000) of the central starburst of J1044+0353 from the COS LegAcy Spectroscopic SurveY (CLASSY, CLASSY; CLASSY2, GO 15840), which allows for detailed modeling of outflow properties along the line of sight without significant effects from the LSF even after downsampling to the 0.073 Å resolution (see CLASSY; CLASSY2; Parker2024). The C ii $\lambda 1334$ absorption line is one of the strongest low ionization lines in the FUV due to the transition oscillator strength and comparatively high ionic abundance, making it an excellent tracer of outflows. Visual inspection of the C ii line profile suggests a classic blue-shifted continuous outflow plus a static ISM component (see, e.g., Figure 1 of 2015ApJ...809..147H). To obtain the normalized flux, we fit the starlight continuum within 3 Å of the feature with a third-order polynomial and divide the local spectrum by this fit (see, e.g., the VoigtFit software, VoigtFit, which takes a similar approach using Chebyshev polynomials).
@@
For modeling the C ii line with OutLines, we use atomic data from Morton2003. No infilling is apparent in the observed line profile, suggesting either maximal aperture clipping where pure absorption is an appropriate description of the emergent feature (see Figure 8) or an optically thick wind. For simplicity, we assume the OutLines defaults, although we note that applying more sophisticated geometry or density profiles negligibly impacts the quality of the fit. Indeed, a filled cone model yields best-fit inclinations of $i\to 0^{\circ}$ and opening angles of $\theta_{o}\to 90^{\circ}$, consistent with the filled sphere geometry. Additional consideration of pulse-like density profiles consistently favors a peak location at the base of the outflow with wide dispersion, which combined are consistent with the power law approximation. Best-fit parameters are in Table 1.
@@
Figure 17: MCMC fit of an OutLines.Absorption model to the C ii $\lambda 1334$ absorption line observed by HST/COS G130M (black, CLASSY). Line styles and colors as in Figure 16. Blue line indicates the intrinsic profile convolved with the HST/COS G130M line spread function (LSF).
The best-fit parameters indicate that the starbursting core of J1044+0353 contains a spherically expanding, isothermal wind with a velocity field characteristic of optically thick, radiatively-driven winds (e.g., 1977ApJ...213..737B; Flury2023). However, the relatively low wind velocity ($v_{\infty}=215$ $\rm km~s^{-1}$, comparable to Voigt modeling of the silicon lines, Parker2024, and the “broad” component of nebular lines, Peng2023) and weak $\dot{p}$ and $\dot{E}$ (marginally non-zero) suggests that mechanical feedback from massive stars is relatively weak or even suppressed. Such suppression can occur in cases where catastrophic cooling of swept up dense material prevents acceleration up to more “traditional” velocities closer to 500 $\rm km~s^{-1}$ or more, is known to occur in Green Pea galaxies with young stellar populations (2017ApJ...851L...9J), and could be associated with Lyman continuum (LyC) escape (e.g., 2017ApJ...851L...9J; Jaskot2019). Combined with evidence for recent SNe, the weak feedback inferred from OutLines models supports the bursty star formation scenario for LyC escape suggested by Flury2022b in their assessment of a large sample of star-forming galaxies with LyC detections and later demonstrated by Flury2025a with measurement of stellar populations and ISM conditions in stacks of UV spectra of LyC emitters. In their framework, SNe clear out channels in the ISM via mechanical feedback and a subsequent generation of stars 8-10 Myr later provides ionizing–but not mechanical–feedback to establish the density-bounded conditions necessary for Lyman continuum photons to escape anisotropically. Here, inferring underlying physics and wind properties from a single-aperture spectrum demonstrates the power and effectiveness of OutLines in understanding the role of mechanical feedback in LyC escape.
@@ 4.3 Outflow Formation and Blowout in a Starburst Galaxy
Mrk 1486 is a starburst galaxy exhibiting a tentative He ii $\lambda 4686$ Wolf-Rayet feature (e.g., Izotov1997), metal rich extraplanar biconical outflows beyond the galactic plane (e.g., Duval2016; Cameron2021), and >14>14 Myr stellar populations (Parker2025, although see Peng2025, who find from the same UV spectra that these populations are even older, $\approx 100$ Myr). The archival HST/COS spectrum from RiveraThorsen2015 exhibits the classic blue-shifted asymmetric feature associated with outflows, albeit with at least two troughs present in the outflow feature across several low ionization lines (Si ii $\lambda\lambda 1191,1193,1304$, O i $\lambda 1302$, and Al ii $\lambda 1670$, see RiveraThorsen2015 their Figure 3, Parker2024, their Figure 7). A multi-troughed absorption feature can indicate a ensemble of shells produced by a series of episodes or a system of many star clusters (see Figure 5, also RiveraThorsen2015).
For our analysis, we use the RiveraThorsen2015 spectrum reprocessed by CLASSY; CLASSY2, down-sampled to 0.073 Å resolution and normalized following the polynomial approach in §4.2. To determine the origin of this multi-episode outflow, we can use OutLines to model the absorption lines. The Si ii $\lambda\lambda 1191,1193$ doublet provides an excellent probe as both features arise from the same ion, thus providing additional constraining power. The stronger Si ii $\lambda 1260$ line is contaminated by telluric O i $\lambda 1302$ and cannot be used to constrain the fit. We note the presence of weak Si ii $\lambda 1194,1197$ fluorescent emission arising from the re-emission of Si ii $\lambda\lambda 1191,1193$ photon absorption, respectively. These lines suffer from apparent Milky Way contamination in the blue wings due to absorption by Galactic O iv $\lambda\lambda 1234,1237$, Si iii $\lambda 1235$, and Mn ii $\lambda\lambda 1234,1236$. We mask these features in order to include the fluorescent lines in our fit. Using atomic data from Kelleher2008 obtained from NIST (NIST_ASD) and the appropriate COS G130M LSF, we model the Si ii doublet complex of $\lambda\lambda 1191,1193$ resonant absorption and emission and $\lambda\lambda 1194,1197$ fluorescent emission using OutLines assuming a damped-pulses density profile to account for the multiple absorption troughs and otherwise adopting the default options of a spherical geometry and the $\beta$ CAK velocity field. The choice of spherical geometry is supported by the presence of relatively symmetric fluorescent Si ii $\lambda 1194,1197$ lines, although the COS LSF may complicate any inference of the geometry. As in §4.2, we incorporate the appropriate LSF corrections.
@@
where $C_{f}$ expresses the fraction of background light which is not obstructed by the foreground gas (e.g., Netzer1982; Arav1999; Arav2005; RiveraThorsen2015). The COS LSF can affect the partial covering fractions of saturated absorption lines (see, e.g., assessment by Flury2025a, their Appendix D, also Figure 3 of Jennings2025). Incorporating both $C_{f}$ and $\epsilon$ as free parameters into the initial fit, we find that while the COS LSF and infilling by the $1194$ fluorescent line can account for some of the line profile depths, $C_{f}=0.82\pm 0.03$ is necessary to provide an optimal description to the absorption features (comparable to the $C_{f}=0.8$ obtained by RiveraThorsen2015). We also find $\epsilon=0.21\pm 0.03$ necessary to describe the fluorescent line profile.
We adjust the prior to require the column densities of each line component to agree within 0.05 dex standard deviation of the column densities of the same components in the other line, effectively enforcing the atomic data. Based on visual assessment of the line profiles, we further impose a Gaussian prior requiring the terminal velocity to agree with the apparent velocity at which the blue profile wing reaches the continuum ($\approx 400$ $\rm km~s^{-1}$, accounting for the LSF) and the central velocity of the first pulse to agree with the velocity of the deepest trough ($\approx 130$ $\rm km~s^{-1}$), both to within a standard deviation of four resolution elements. We also impose that each shell is relatively thin by requiring the width of each shell to be $15\pm$ 15% of the launch radius. We show the best-fit line profile in Figure 18 and list the results in Table 1. As a test of our assumption of spherical geometry, we predict the emergent fluorescent Si ii $\lambda\lambda 1194,1197$ lines using OutLines.Fluorescent and the best-fit parameters from the posterior, which we show in Figure 18.
@@
Figure 18: MCMC fit of an OutLines model to the Si ii $\lambda\lambda 1191,1193,1194,1997$ doublet complex observed by HST/COS G130M (black, RiveraThorsen2015, GO 12583, PI Hayes). Shaded regions indicate suspected Milkyway contamination from Galactic Si iii, Mn $i$, and and O iv. Si ii $\lambda\lambda 1191,1193$ resonant emission lines computed using OutLines.Resonant are shown in orange, demonstrating strong radiative trapping. Si ii $\lambda\lambda 1194,1197$ fluorescent lines computed using OutLines.Fluorescent are shown in yellow. Total line profile corrected for the line spread function (LSF) is shown in blue. Other line styles, symbols, and colors as in Figure 16.
@@
The posterior sampling indicates relatively small uncertainties on the order of 10% for each of the expanding shells, with the most prominent and lowest velocity shell having $v_{\rm shell}=188\pm 20$ $\rm km~s^{-1}$ (obtained from $v_{\rm shell}=v_{\infty}w[x_{1}]$ and propagating the uncertainties in $x_{1}$, $\beta$, and $v_{\infty}$). The system of shells suggest an outflow in formation, with regular events forming thin expanding superbubbles. Notably, the less pronounced absorption troughs in C ii $\lambda 1334$, which traces a higher ionization state than Si ii, O i, and Al ii, suggests that the ionized phase of the outflow is established more rapidly than the neutral phase and may indicate that the expanding bubbles are entrained in or even driven by a hotter phase of the ISM. The accumulation of superbubbles close to the terminal velocity suggests that an outflow forms from these successive events with sufficient speed to exceed the escape velocity ($v_{esc}\approx$100 $\rm km~s^{-1}$as implied by the $M_{\star}=10^{9.3}\rm~M_{\odot}$ from 2018MNRAS.481.1690C). The recurrance of many such expanding shells could produce the biconical outflow observed in emission once the superbubbles escape the gravitational potential and blow out from the galactic disk (see spatially resolved cases in, e.g., Martin1998; McQuinn2019, and Duval2016 specifically regarding Mrk 1486). Moreover, the successive events are removing gas at mass outflow rates of $\dot{M}=7.9\rm~M_{\odot}~yr^{-1}$, which may temporarily rob the galaxy of baryonic material necessary to sustain the production of stars and could explain the 30-40 Myr lapse between bursts of star formation.
@@
Given that the outflows in Mrk 1486 are oxygen rich relative to their host galaxy (Cameron2021), it is useful to consider whether the outflow component of Si ii is metal-enhanced since, like oxygen, silicon is an $\alpha$ element and is predominantly produced by core-collapse supernovae (CCSNe, e.g., Kobayashi2020). From the best-fit line profiles from OutLines, the outflow $N_{o\ \rm Si^{+}}=10^{15.5}\rm~cm^{-2}$ while the static $N_{s\ \rm Si^{+}}=10^{13.5}\rm~cm^{-2}$. Ionization structure may account for some of the difference in Si+; nevertheless, the 2 dex increase in the outflow likely suggests substantial metal loading of the outflow (see, e.g., 2018MNRAS.481.1690C) due to enrichment by recent CCSNe.

### v2

@@ 4.1 Evolution of H ii Regions
H ii regions are dynamic environments with radiation and winds from O and B stars shaping and enriching the proximal ISM. Winds launched from stars carry out recently formed metals dredged up to the surface and subsequently sweep up surrounding material. Radiation heats gas, often through photoionization, and imparts momentum to the ISM, which can result in additional launching or driving of winds, bubbles, and outflowing gas (e.g., Castor et al., 1975b; Weaver et al., 1977; Dyson, 1977, 1979; Oey and Clarke, 1997; Dopita et al., 2005). While stellar population models often contain predictions for these feedback effects (e.g., Leitherer et al., 1999, 2014; Eldridge et al., 2017; Byrne et al., 2022; Hawcroft et al., 2025), empirically constraining the kinematics and related mechanical feedback in H ii regions remains an active area of research (e.g., models and simulations by Geen et al., 2020, 2023; Lancaster et al., 2025b, a) with implications for understanding chemical enrichment (e.g., Esteban and Peimbert, 1995; López-Sánchez et al., 2007; Stock et al., 2011) and dispersal timescales (e.g., Dopita et al., 2005; Menon et al., 2024).
One such kinematically rich and complex system is the giant extragalactic H ii region NGC 5471 located in M101. Many knots within this region exhibit broad and sometimes asymmetric nebular lines which cannot be described by a single Gaussian (e.g., Chu and Kennicutt, 1986; Bresolin et al., 2020). To investigate the evolution of H ii regions in this case-study, we apply OutLines to the [S ii] $\lambda\lambda 6716,6731$ doublet, which traces the density and kinematics of low ionization gas, in Knot A of NGC 5471, the brightest knot identified by Skillman (1985), as observed with the $R=6000$ MEGARA IFU spectrograph on the Gran Telescopia Canarias (Arellano-Cordova in prep, PI Esteban). For our modeling, we use a 2.5′′ aperture extraction of the Knot A spectrum from the IFU cube. We obtain atomic data from NIST (Kramida et al., 2024) for wavelengths (Kaufman and Martin, 1993) and spontaneous emission coefficients (transition probabilities) (Podobedova et al., 2009). We use collision strengths from Tayal and Zatsarinny (2010). First, we fit the adjacent featureless continuum with a third-order polynomial, an appropriate approximation in the case of extended H ii regions which do not suffer appreciably from contamination such as stellar absorption except near permitted lines from, e.g., H i and He i. We subtract this continuum before modeling the line profile with OutLines with the default options. Then, we sample the posterior on the OutLines profile parameters to constrain the outflow and static gas properties. No LSF corrections are necessary as $R=6000$ is sufficiently high to constrain the kinematics (see §3.2). We compare the best-fit model to the observed line profiles in Figure 17.
@@
From the best-fit line fluxes, we determine the electron density of both the nebula and outflow components from the [S ii] 6716/6731{6716}/{6731} flux ratio using PyNeb (Luridiana et al., 2015). The static gas flux ratio of 1.459 is in the low-density limit, requiring $n_{e}<10\rm\ cm^{-3}$. For the outflowing gas component, the doublet ratio is 0.831, which is where [S ii] is most sensitive simultaneously to density and to temperature. For a range in $T_{e}$ from 10 to 15 $\rm kK$, the inferred outflow density is anywhere from $970$ to $1140\rm~cm^{-3}$, with density values increasing with the assumed temperature. The implication from these densities is that the outflow is roughly two orders of magnitude more dense than the gas within the static nebular gas. While this increase in density may be due to a sweeping up of material in the wind, this dramatic change in density coupled with the supersonic flow implied by $v_{\rm out}$ could instead arise from compression by shocked gas as it rapidly cools via radiation (e.g., Flury et al., 2025a).
Though the posterior on $\beta$ is not well constrained, the slope of the velocity field strongly favors $\beta>4$, indicating gradual acceleration likely caused by optically thick radiation pressure from early type stars (e.g., Barlow and Cohen, 1977). While Bresolin et al. (2020) argue that the non-thermal radio SEDs of Sramek and Weedman (1986) are clear-cut evidence for supernova-driven bubbles, the line profile shape indicates a continuous wind rather than shells or bubbles and, from modeling with OutLines, disfavors the supernova scenario. Indeed, the OutLines results are very much in agreement with the fact that Knot A features the a very young ($3\rm~Myr$) cluster–in fact the youngest in the entirety of NGC 5471 (e.g., García-Benito et al., 2011, based on HST photometric color-magnitude diagrams). Knot A is also very massive ($10^{4}\rm~M_{\odot}$) and requires a significant population of O stars to produce the $Q=4\times 10^{51}\rm~phot~s^{-1}$ associated with the star cluster (García-Benito et al., 2011). The age and mass of the stellar population indicate sufficient radiation available in the UV to radiatively drive a supersonic wind. Shocks triggered by this wind could explain the non-thermal emission in the radio in addition to the dense gas. Inhomogeneities in the gas distribution can readily dissipate the outward force due to pressure from the hot shocked gas (e.g., Dyson, 1977), allowing radiation pressure to dominate the acceleration of material out from the H ii region. Our results indicating escape of hot gas are consistent with the Harper-Clark and Murray (2009) model for the Carina Nebula and the Lopez et al. (2011) multiwavelength study of the massive extragalactic H ii region 30 Doradus, both of which found strong evidence for radiation pressure driving and a deficit of X-ray flux associated with shocked gas.
@@ 4.2 Lyman Continuum Escape from a Super Star Cluster
Starburst galaxy J1044+0353 is a low mass ($M_{\star}\approx 10^{6.8}\rm~M_{\odot}$) Lyman continuum emitter candidate with $f_{esc}^{\rm Lyc}$ as high as 60% (Parker et al., 2025) and, as a nearby ($z=0.01287$) Green Pea, is a local analog for star-forming galaxies at cosmic dawn that are actively involved in re-ionizing the Universe (Martin et al., 2024). Recent IFU observations of J1044+0353 reveal expanding shells of gas at the periphery of the nebular gas flux distribution (Martin et al., 2024), indicative of blowout due to SNe from an aging stellar population 10-20 Myr after a burst of star formation (Peng et al., 2023). However, the central core of the galaxy hosts a super star cluster with several young (<2<2 Myr) stellar populations which may be responsible for driving additional feedback on shorter timescales (Olivier et al., 2022), providing a unique laboratory to explore the impact of young stars on their immediate environment in the context of stochastic, or bursty, star formation using OutLines.
@@
To undertake this investigation, we draw on archival HST/COS G130M FUV spectroscopy (R=10  [^0] R=10\ 000) of the central starburst of J1044+0353 from the COS LegAcy Spectroscopic SurveY (CLASSY, Berg et al., 2022; James et al., 2022, GO 15840), which allows for detailed modeling of outflow properties along the line of sight without significant effects from the LSF even after downsampling to the 0.073 Å resolution (see Berg et al., 2022; James et al., 2022; Parker et al., 2024). The C ii $\lambda 1334$ absorption line is one of the strongest low ionization lines in the FUV due to the transition oscillator strength and comparatively high ionic abundance, making it an excellent tracer of outflows. Visual inspection of the C ii line profile suggests a classic blue-shifted continuous outflow plus a static ISM component (see, e.g., Figure 1 of Heckman et al., 2015). To obtain the normalized flux, we fit the starlight continuum within 3 Å of the feature with a third-order polynomial and divide the local spectrum by this fit (see, e.g., the VoigtFit software, Krogager, 2018, which takes a similar approach using Chebyshev polynomials).
@@
For modeling the C ii line with OutLines, we use atomic data from Morton (2003). No infilling is apparent in the observed line profile, suggesting either maximal aperture clipping where pure absorption is an appropriate description of the emergent feature (see Figure 8) or an optically thick wind. For simplicity, we assume the OutLines default geometry and density profile, although we note that applying more sophisticated geometry or density profiles negligibly impacts the quality of the fit. Indeed, a filled cone model yields best-fit inclinations of $i\to 0^{\circ}$ and opening angles of $\theta_{o}\to 90^{\circ}$, consistent with the filled sphere geometry. Additional consideration of pulse-like density profiles consistently favors a peak location at the base of the outflow with wide dispersion, which combined are consistent with the power law approximation. However, the sudden drop in flux immediately blueward of the systemic velocity of the system suggests a non-zero launch velocity. We therefore allow for non-zero launch velocities. The OutLines default top hat (uniform-bounded) prior is assumed. Best-fit parameters are in Table 1.
@@
Figure 18: MCMC fit of an OutLines.Absorption model to the C ii $\lambda 1334$ absorption line observed by HST/COS G130M (black, Berg et al. 2022). Line styles and colors as in Figure 17. Blue line indicates the intrinsic profile convolved with the HST/COS G130M line spread function (LSF).
The best-fit parameters indicate that the starbursting core of J1044+0353 contains a spherical expanding wind with a velocity field characteristic of radiatively-driven winds most likely due to dusty radiation pressure (e.g., Flury et al., 2023), although line-driving (Castor et al., 1975a) and optically thick (Barlow and Cohen, 1977) mechanisms are consistent with the range of $\beta$. Critically, the posterior on $\beta$ strongly disfavors ram pressure driving, suggesting that supernovae do not cause the current feedback. However, the relatively low wind velocity ($v_{\infty}=215$ $\rm km~s^{-1}$, comparable to Voigt modeling of the silicon lines, Parker et al. 2024, and the “broad” component of nebular lines, Peng et al. 2023) and weak $\dot{p}$ and $\dot{E}$ (marginally non-zero) suggests that mechanical feedback from massive stars is relatively weak or even suppressed. Such suppression can occur in cases where catastrophic cooling of swept up dense material prevents acceleration up to more “traditional” velocities closer to 500 $\rm km~s^{-1}$ or more, is known to occur in Green Pea galaxies with young stellar populations (Jaskot et al., 2017), and could be associated with Lyman continuum (LyC) escape (e.g., Jaskot et al., 2017, 2019). Combined with evidence for recent SNe, the weak feedback inferred from OutLines models supports the bursty star formation scenario for LyC escape suggested by Flury et al. (2022) in their assessment of a large sample of star-forming galaxies with LyC detections and later demonstrated by Flury et al. (2025b) with measurement of stellar populations and ISM conditions in stacks of UV spectra of LyC emitters. In their framework, SNe clear out channels in the ISM via mechanical feedback and a subsequent generation of stars 8-10 Myr later provides ionizing–but not mechanical–feedback to establish the density-bounded conditions necessary for Lyman continuum photons to escape anisotropically. Here, inferring underlying physics and wind properties from a single-aperture spectrum demonstrates the power and effectiveness of OutLines in understanding the role of mechanical feedback in LyC escape.
@@ 4.3 Outflow Formation and Blowout in a Starburst Galaxy
Mrk 1486 is a starburst galaxy exhibiting a tentative He ii $\lambda 4686$ Wolf-Rayet feature (e.g., Izotov et al., 1997), metal rich extraplanar biconical outflows beyond the galactic plane (e.g., Duval et al., 2016; Cameron et al., 2021), and >14>14 Myr stellar populations (Parker et al. 2025, although see Peng et al. 2025, who find from the same UV spectra that these populations are even older, $\approx 100$ Myr). The archival HST/COS spectrum from Rivera-Thorsen et al. (2015, GO 12583, PI Hayes) exhibits the classic blue-shifted asymmetric feature associated with outflows, albeit with at least two troughs present in the outflow feature across several low ionization lines (Si ii $\lambda\lambda 1191,1193,1304$, O i $\lambda 1302$, and Al ii $\lambda 1670$, see Rivera-Thorsen et al. 2015 their Figure 3, Parker et al., 2024, their Figure 7). A multi-troughed absorption feature can indicate a ensemble of shells produced by a series of episodes or a system of many star clusters (see Figure 5, also Rivera-Thorsen et al. 2015).
For our analysis, we use the Rivera-Thorsen et al. (2015) spectrum reprocessed by Berg et al. (2022); James et al. (2022), down-sampled to 0.073 Å resolution and normalized following the polynomial approach in §4.2. To determine the origin of this multi-episode outflow, we can use OutLines to model the absorption lines. The Si ii $\lambda\lambda 1191,1193$ doublet provides an excellent probe as both features arise from the same ion, thus providing additional constraining power. The stronger Si ii $\lambda 1260$ line is contaminated by telluric O i $\lambda 1302$ and cannot be used to constrain the fit. We note the presence of weak Si ii $\lambda 1194,1197$ fluorescent emission arising from the re-emission of Si ii $\lambda\lambda 1191,1193$ photon absorption, respectively. These lines suffer from apparent Milky Way contamination in the blue wings due to absorption by Galactic O iv $\lambda\lambda 1234,1237$, Si iii $\lambda 1235$, and Mn ii $\lambda\lambda 1234,1236$. We mask these features in order to include the fluorescent lines in our fit. Using atomic data from Kelleher and Podobedova (2008) obtained from NIST (Kramida et al., 2024) and the appropriate COS G130M LSF, we model the Si ii doublet complex of $\lambda\lambda 1191,1193$ resonant absorption and emission and $\lambda\lambda 1194,1197$ fluorescent emission using OutLines assuming a damped-pulses density profile to account for the multiple absorption troughs and otherwise adopting the default options of a spherical geometry and the $\beta$ CAK velocity field. The choice of spherical geometry is supported by the presence of relatively symmetric fluorescent Si ii $\lambda 1194,1197$ lines, although the COS LSF may complicate any inference of the geometry. As in §4.2, we incorporate the appropriate LSF corrections.
@@
where $C_{f}$ expresses the fraction of background light which is not obstructed by the foreground gas (e.g., Netzer, 1982; Arav et al., 1999, 2005; Rivera-Thorsen et al., 2015). The COS LSF can affect the partial covering fractions of saturated absorption lines (see, e.g., assessment by Flury et al., 2025b, their Appendix D, also Figure 3 of Jennings et al. 2025). Incorporating both $C_{f}$ and $\epsilon$ as free parameters into the initial fit, we find that while the COS LSF and infilling by the $1194$ fluorescent line can account for some of the line profile depths, $C_{f}=0.82\pm 0.03$ is necessary to provide an optimal description to the absorption features (comparable to the $C_{f}=0.8$ obtained by Rivera-Thorsen et al., 2015). We also find $\epsilon=0.21\pm 0.03$ necessary to describe the fluorescent line profile.
We show the best-fit line profile in Figure 19 and list the results in Table 1. As a test of our assumption of spherical geometry, we predict the emergent fluorescent Si ii $\lambda\lambda 1194,1197$ lines using OutLines.Fluorescent and the best-fit parameters from the posterior, which we show in Figure 19.
@@
Figure 19: MCMC fit of an OutLines model to the Si ii $\lambda\lambda 1191,1193,1194,1997$ doublet complex observed by HST/COS G130M (black, Rivera-Thorsen et al. 2015, GO 12583, PI Hayes). Shaded regions indicate suspected Milkyway contamination from Galactic Si iii, Mn $i$, and and O iv. Si ii $\lambda\lambda 1191,1193$ resonant emission lines computed using OutLines.Resonant are shown in orange, demonstrating strong radiative trapping. Si ii $\lambda\lambda 1194,1197$ fluorescent lines computed using OutLines.Fluorescent are shown in yellow. Total line profile corrected for the line spread function (LSF) is shown in blue. Other line styles, symbols, and colors as in Figure 17.
@@
The posterior sampling indicates relatively small uncertainties on the order of 10% for each of the expanding shells, with the most prominent and lowest velocity shell having $v_{\rm shell}=188\pm 20$ $\rm km~s^{-1}$ (obtained from $v_{\rm shell}=v_{\infty}w[x_{1}]$ and propagating the uncertainties in $x_{1}$, $\beta$, and $v_{\infty}$). The system of shells suggest an outflow in formation, with regular events forming thin expanding superbubbles. Notably, the less pronounced absorption troughs in C ii $\lambda 1334$, which traces a higher ionization state than Si ii, O i, and Al ii, suggests that the ionized phase of the outflow is established more rapidly than the neutral phase and may indicate that the expanding bubbles are entrained in or even driven by a hotter phase of the ISM. The accumulation of superbubbles close to the terminal velocity suggests that an outflow forms from these successive events with sufficient speed to exceed the escape velocity ($v_{esc}\approx$100 $\rm km~s^{-1}$as implied by the $M_{\star}=10^{9.3}\rm~M_{\odot}$ from Chisholm et al. 2018). The recurrance of many such expanding shells could produce the biconical outflow observed in emission once the superbubbles escape the gravitational potential and blow out from the galactic disk (see spatially resolved cases in, e.g., Martin, 1998; McQuinn et al., 2019, and Duval et al. 2016 specifically regarding Mrk 1486). Moreover, the successive events are removing gas at mass outflow rates of $\dot{M}=4.8\rm~M_{\odot}~yr^{-1}$, which may temporarily rob the galaxy of baryonic material necessary to sustain the production of stars and could explain the 30-40 Myr lapse between bursts of star formation.
@@
Given that the outflows in Mrk 1486 are oxygen rich relative to their host galaxy (Cameron et al., 2021), it is useful to consider whether the outflow component of Si ii is metal-enhanced since, like oxygen, silicon is an $\alpha$ element and is predominantly produced by core-collapse supernovae (CCSNe, e.g., Kobayashi et al., 2020). From the best-fit line profiles from OutLines, the outflow $N_{o\ \rm Si^{+}}=10^{15.5}\rm~cm^{-2}$ while the static $N_{s\ \rm Si^{+}}=10^{13.5}\rm~cm^{-2}$. Ionization structure may account for some of the difference in Si+; nevertheless, the 2 dex increase in the outflow likely suggests substantial metal loading of the outflow (see, e.g., Chisholm et al., 2018) due to enrichment by recent CCSNe.